merah13:

What is the solution to the equation StartFraction y Over y minus 4 EndFraction minus StartFraction 4 Over y + 4 EndFraction = StartFraction 32 Over y squared minus 16 EndFraction? y = –4 and y = 4 y = 0 all real numbers no solution

Vocaloid:

\[\frac{ y }{ y-4 }-\frac{ 4 }{ y+4 }=\frac{ 32 }{ y^{2}-16 }\] y^2 - 16 is a difference of squares. so it can be factored into (y+4)(x-4) \[\frac{ y }{ y-4 }-\frac{ 4 }{ y+4 }=\frac{ 32 }{ (y+4)(y-4)}\] now, notice how the third denominator has the two factors (y+4) and (y-4). each denominator on the left side has one of these factors. for each left fraction, multiply by the appropriate factor to convert the denominator to (y+4)(y-4). so for y / (y-4), it's missing a y + 4 in the denominator, so we multiply the whole fraction by (y+4)/(y+4). \[\frac{ y(y+4) }{ (y+4)(y-4) }-\frac{ 4 }{ y+4 }=\frac{ 32 }{ (y+4)(y-4) }\] same logic with the second fraction. multiply by (y-4)/(y-4) \[\frac{ y(y+4) }{ (y+4)(y-4) }-\frac{ 4(y-4) }{ (y+4)(y-4) }=\frac{ 32 }{ (y+4)(y-4) }\] now that all the denominators are equal, you can eliminate the denominators by multiplying both sides by (y+4)(y-4) y(y+4) - 4(y-4) = 32 finish by distributing, combining like terms, and isolating y